OpenStudy (anonymous):

1.7x + 8 - 1.62x = 0.4 x - 0.32 + 8

OpenStudy (anonymous):

x=-(0,32/0,4)

OpenStudy (anonymous):

or -0.8

OpenStudy (anonymous):

1.7x-1.62x=.08x therefore, .08x+8=.4x+7.68 8-7.68=-.08x+.4x .32=.32x 1=x.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

u are wrong. -.08x+.04x makes 0.4x not 0.32